OpenStudy (anonymous):

Select the property that allows the left member of the equation to be changed to the right member. 8 + 0 = 8

OpenStudy (mathstudent55):

When 0 is added to any number, the number remains the same. For example, 5 + 0 = 5 8 + 0 = 8 In general, for any number a, a + 0 = a. You could say that the "identity" of a number does not change by adding zero to it. That helps remember the name of the property. The "additive identity property."

OpenStudy (anonymous):

thankyou

OpenStudy (mathstudent55):

You're welcome.